OpenStudy (anonymous):

(super easy question) what percent of 230 is 23 (just explain i forgot again)

OpenStudy (anonymous):

Well, think about it. 230 times what is 23?

OpenStudy (anonymous):

230x = 23, x = .1

OpenStudy (anonymous):

.1 is ten percent, so that's your answer.

OpenStudy (calculusxy):

% of 230=23 is % -- = --- of 100 is=23 of=230 %= x (what we are looking for) 23 1 -- simplified -- 230 10 1 x -- = ---- 10 100 solve for x: 100=10x x=10

OpenStudy (anonymous):

23/230=1/10=10%